OpenStudy (anonymous):

Have any questions??? I can help! I can not help all questions but I'll try to! Don't understand something come to me!

OpenStudy (anonymous):

Compare and Contrast. Below are two sets of points. Find the slopes between the two points and compare them. Choose the statement that is true about the values of the slopes. Set #1 Set #2 (4, 4) and (2, 6) (2, 4) and (4, 2) The slope of Set #1 is smaller than the slope of Set #2. The slope of Set #1 is larger than the slope of Set #2. The slope of Set #1 is the same as the slope of Set #2. None of the statements above describes slopes of sets shown.

OpenStudy (anonymous):

solve algebraically for all value of theta in the interval 0 degrees less than or equal to theta and less than or equal to 360 degrees. 2sin^2 theta-4sin theta=cos^@ theta-2. express in degrees

OpenStudy (anonymous):

A principal of $835 is invested in an account at 2 percent per quarter simple interest. which of the following sequences describes the dollar amount of principal amount at quarterly intervals? (there can be more than one correct answer) A. 66.80, 133.60, 200.40, 267.20, 334.00,... B. 835.00, 851.70, 868.73, 886.11, 903.83,... C. 16.70, 33.40, 50.10, 66.80, 83.50,... D. 835.00, 851.70, 868.40, 885.10, 901.80,... E. 835.00, 901.80, 968.60, 1035.40, 1102.20,...

OpenStudy (anonymous):

@brittbrat24 Do you know the equation for computing a slope?

OpenStudy (anonymous):

@mangorox no can you tell me the answer ?

OpenStudy (anonymous):

Ok, slope is rise divided by the run. So, do that for set 1 and then set 2 and compare them and that should give you the answer.

OpenStudy (anonymous):

I'll give you an example of rise and run hold on

OpenStudy (anonymous):

solve algebraically for all value of theta in the interval 0 degrees less than or equal to theta and less than or equal to 360 degrees. 2sin^2 theta-4sin theta=cos^@ theta-2. express in degrees

OpenStudy (anonymous):

 @brittbrat24

OpenStudy (anonymous):

@mangorox so the answer would be B right ?

OpenStudy (anonymous):

@brittbrat24 Exactly! The answer is B! Great job!

OpenStudy (anonymous):

@mangorox thank you !!!! :)

OpenStudy (anonymous):

@brittbrat24 Anytime!

OpenStudy (anonymous):

What is the slope of the line 2x + 5y = -10?

OpenStudy (anonymous):

Think of it this way: 2*? + 5*? = -10

OpenStudy (anonymous):

I don't really know how to explain that question otherwise sorry!

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Which one of these equations is correctly written in Standard Form? x + y = −4 3x + 2.5y = −4 x + 3y = −12 −3x + y = −12

OpenStudy (anonymous):

@brittbrat24 Standard form: Ax + By = C, where A, B, C are integers with greatest common factor = 1, and A > 0 Answer: c. x + 3y = −12

OpenStudy (anonymous):

@mangorox yae !!!!! i had the same same answer as you

OpenStudy (anonymous):

@brittbrat24 Great!

OpenStudy (anonymous):

To graph the equation y= -1^2x-3 using the slope and y-intercept, you would have to do which of the following? Graph the y-intercept, 3. From that point, move down 1 and to the right 2 to plot another point. Then draw a line that passes through both points. Graph the y-intercept, -3. From that point, move up 1 and to the right 2 to plot another point. Then draw a line that passes through both points. Graph the y-intercept, -3. From that point, move down 1 and to the right 2 to plot another point. Then draw a line that passes through both points. Graph the y-intercept, 3. From that point, move up 1 and to the right 2 to plot another point. Then draw a line that passes through both points.

OpenStudy (anonymous):

I honestly don't know this one sorry :( @brittbrat24

OpenStudy (anonymous):

-lol ok i have another one

OpenStudy (anonymous):

Using complete sentences, explain how to graph y = 1^3x + 4 using the slope-intercept method.

OpenStudy (anonymous):

so your equation is y=1/3x + 4? alright let me break this equation down. whatever is multiplied by x is the slope (aka rate of change) so slope= 1/3 and ill come back to that. whatever is being added or subtracted to the x is where the line starts. so since its +4 that means the graph starts at 4 on the Y axis.. so 4 points up. (example: if it was y= 1/3x -3 the line would start down 3 on the y axis) so the y intercept is 4 because the line will be starting on/crossing the y axis at 4 back to the slope! slope can be remembered as rise over run (aka rise/run where 1=rise and 3=run because the slope is 1/3 in this equation). im sure you've heard this before, but it means that from the starting point (which we found is 4 on the y axis) we will rise 1 up and run 3 to the right and plot a new point in that spot. then repeat this step. rise 1 up and run 3 over. by this time you will have 3 points and you dont really need more than 2. Connect them and there is your finished line! so using complete sentences.. just explain that 1/3 is the slope and you start 4 up on the y axis and you draw your line by using rise over run with the slope.. but word it yourself. BTW: slope-intercept method is just using the equation " y= mx + b " where m= slope (1/3 in this case) and b = the starting point on the y axis (4 in this case)

OpenStudy (anonymous):

Simpler/shorter version: Hmmm ...not if we are using the slope intercept method ... I think after we plot (0,4) then we use the slope by rising 1 and running three to the right, this will land us at (3, 4). Now we connect the two points. @brittbrat24

OpenStudy (anonymous):

@mangorox thank you

OpenStudy (anonymous):

@brittbrat24 No problem

OpenStudy (anonymous):

A principal of $835 is invested in an account at 2 percent per quarter simple interest. which of the following sequences describes the dollar amount of principal amount at quarterly intervals? (there can be more than one correct answer) A. 66.80, 133.60, 200.40, 267.20, 334.00,... B. 835.00, 851.70, 868.73, 886.11, 903.83,... C. 16.70, 33.40, 50.10, 66.80, 83.50,... D. 835.00, 851.70, 868.40, 885.10, 901.80,... E. 835.00, 901.80, 968.60, 1035.40, 1102.20,...